the 10th term of an AP is -27 and the 5th term is -12. what is the 18th term find also the sum of its 28th terms
Answer by stanbon(75887)

the 10th term of an AP is -27 and the 5th term is -12. what is the 18th term find also the sum of its 28th terms
------
a(10) = a(1) + 9d = -27
a(5) = a(1) + 4d = -12
------------------------------
Subtract and solve for "d"::
5d = -13
d = -13/5
------------
Solve for a(1):::
a(1) + 4d = -12
a(1) -52/5 = -12
a(1) = -60/5 + 52/5
a(1) = -8/5
----------------------
Ans: a(18) = a(1)+17d = -8/5 + 17(-13/5) = -8/5 - 221/5 = -229/5 = -15 4/5
----------------
To get the sum of the 28 terms
S(28) = (28/2)(a(1)+a(28))
====
I'll leave that to you.
Cheers,
Stan H.
----------